# Online Multi-Task Offloading for Semantic-Aware Edge Computing Systems

### v2

@@ Abstract
Index Terms:
Mobile edge computing, resource allocation, task offloading, semantic-aware.
@@ I Introduction
The sixth generation (6G) of mobile communication is demonstrating its revolutionary potential, portending the arrival of an era of ultra-massive connectivity and the Internet of Things (IoT). The volume of data generated by a vast number of devices is experiencing an explosive surge. Various compute-intensive services, such as holographic communication, extended reality (XR), and autonomous driving continue to evolve, which increasingly burden user equipments (UEs) with heavy computational demands. However, the limited computational power and battery capacity of UEs make it challenging to support a multitude of computationally intensive tasks. Cloud computing is considered capable of alleviating the computational pressure on UEs to some extent [1]. Offloading tasks to the cloud, which has abundant computing resources, significantly reduces local computational pressure. However, the high latency associated with cloud offloading is intolerable for latency-sensitive tasks. Furthermore, the centralized model of cloud computing poses data privacy and security issues. To address the aforementioned challenges, mobile edge computing (MEC) is considered an effective solution. Edge servers (ES) near the user can provide considerable computing and storage resources, which can be fully utilized by UEs distributed. Compared to cloud computing, MEC significantly reduces latency, while its distributed computing resources ensure user privacy and security.
@@
Nevertheless, when tasks involve substantial data transfers, such as high-definition (HD) video processing [2], virtual reality (VR) applications [3], or real-time data analytics [4], the increased volume of data transmission imposes higher demands on channel quality to meet latency requirements. In such cases, semantic communication (SemCom) can provide an effective solution. SemCom, different from the traditional bit workflow based on Shannon’s information theory [5], has demonstrated great potential for enhancing the efficiency and reliability of data transmission across various task scenarios, including text [6, 7, 8], image [9, 10, 11], video [12, 13, 14], XR [15, 16, 17], and multi-task [18, 19, 20]. By transmitting a compact semantic representation of task data, substantial bandwidth can be saved, ensuring high quality of service (QoS).
I-A Related Works
MEC technology has attracted widespread attention from academia and industry [21] due to its ability to enhance the computing capabilities of UEs and improve the quality of service (QoS). In [22], the authors optimized the allocation of communication, computation, and energy resources based on a perturbed Lyapunov algorithm, maximizing system throughput under the stability constraints of task and energy queues. A power minimization method with task buffer stability constraints was proposed in [23], where the Lyapunov optimization algorithm was used to solve computing offloading strategies. Liu et al. considered the requirements for ultra-reliable and ultra-low latency (URLLC) tasks [24], applying extreme value theory to minimize user power consumption while satisfying probabilistic queue length constraints.
However, the aforementioned model-based approaches are typically computationally intensive, further burdening user devices. Moreover, model-based methods often require multiple iterations to find the optimal offloading strategy, which is intolerable for latency-sensitive tasks and impractical in rapidly changing wireless channel conditions. To address these issues, resource allocation and computation offloading methods based on Deep Reinforcement Learning (DRL) were proposed. The authors of [25] modeled the computation offloading as a multi-stage stochastic mixed-integer nonlinear programming (MINLP) problem and solved it using a combination of Lyapunov optimization and a DRL model, integrating the advantages of both model-based and DRL approaches. In [26, 27], the authors exclusively utilized DRL to solve for joint strategies of computation offloading and resource allocation, similarly achieving quite impressive results.
For services like holographic communication [28] and real-time video analysis [29], which are both latency-sensitive and compute-intensive, ensuring consistent QoS is challenging due to the need to transmit large amounts of data while combating wireless channel fading. Elaborate deep neural networks are often used for such sophisticated tasks, however, they also introduce significant computational and storage burdens on devices, necessitating the leverage of ES computing power through integration with MEC. Although offloading tasks to the ES can alleviate the computational burden on UEs and reduce execution latency, excessive data offloading or too many UEs offloading simultaneously can lead to network congestion, resulting in offloading failures.
@@ I-A Related Works
I-B Motivations and Contributions
In traditional MEC scenarios, computational and communication resources are often jointly optimized to minimize system energy consumption [35] or execution latency [36], or a combination of both [37]. However, these approaches lack consideration for task execution performance and comprehensive QoE. 6G strives to provide stable, reliable, and high-quality personalized services to each user [38]. Preferences often vary among different users and types of services. For instance, in MEC scenarios, users dealing with latency-sensitive tasks may prefer lower latency, those pursuing task performance may favor better task execution accuracy, while users with limited battery power might opt for task offloading strategies that consume less power. Therefore, a good QoE metric should encompass the preferences of different users. Additionally, since each user may need to handle different tasks, the complexity and evaluation metrics of different tasks vary. For example, bimodal tasks like visual question answering (VQA) typically require more processing time and computational resources compared to text translation, and their evaluation metrics are entirely different.
Semantic-aware MEC systems have been studied in [31, 32, 33, 34]. Nonetheless, in [31, 32, 33], the authors did not consider multi-tasks, limiting their ability to compare users engaged in different types of tasks. The authors of [34] did not consider the impact of semantic extraction on task execution latency and performance. To address the aforementioned challenges, we investigate resource allocation schemes in multi-task offloading that coordinate computational and communication resources through semantic compression-aware methods. To fairly compare the offloading performance of different task types, a unified QoE metric is proposed. Computational and communication resources are jointly optimized to maximize the overall QoE for UEs. The main contributions of this paper are summarized as follows:
  • We propose a novel semantic-aware task offloading system for multi-tasks. The SemCom systems for image classification, text classification, and VQA are first constructed to compress source data into compact semantic representations. The semantic extraction factor is used to characterize the degree of data compression, which affects the task’s execution latency, energy consumption, and performance.
@@ I-B Motivations and Contributions
The remainder of this paper is organized as follows. Section II describes the semantic-aware transmission and computation model, defines the unified QoE metric, and formulates the problem as maximizing overall QoE. Section III introduces solving the QoE maximization problem using the MAPPO algorithm, providing training details. Section IV presents relevant experimental details and simulation results. Finally, Section V outlines our conclusions.
II System Model and Problem Formulation
@@
We propose a novel semantic-aware multi-task offloading system, as shown in Fig. 1, where UEs employ an orthogonal frequency-division multiple access (OFDMA) scheme. At the access point (AP), $K$ resource blocks are allocated to serve $N$ users at different sub-channels. The set of UEs and sub-channels are defined as $\mathcal{N}=\{1,2,\ldots,N\}$ and $\mathcal{K}=\{1,2,\ldots,K\}$. The length of the UEn’s task queue is denoted as $Q_{n}$. The task currently being executed by the UEn is represented by $q^{\mathcal{M}}_{n}$, where $\mathcal{M}\in\{task_{1},task_{2},\ldots,task_{m}\}$. Hence, each UE may contain up to $m$ different types of tasks. We assume that tasks are atomic [34], i.e., all tasks are either executed locally or offloaded to the ES. The illustrations of the main notations are summarized in Table I.
II-A Semantic-Aware Model
@@
Traditional MEC offloading methods usually require transmitting all data to the ES. SemCom enables the extraction of task-relevant semantic information, reducing bandwidth consumption while enhancing robustness to channel variations. Based on this, we propose a semantic-aware multi-task offloading framework. Fig. 2 illustrates the process of semantic-aware offloading. While semantic extraction indeed reduces the volume of data transmitted, the associated increase in semantic extraction latency and computational burden must also be comprehensively considered. Hence, we define a semantic extraction factor $\mu_{n}\in\left[\mu_{n}^{min},1\right]$. The semantic extraction module determines the level of data compression. When $\mu_{n}=1$, the UE transmits all data instead of semantic information. Transmitting all data can achieve better task performance, but it imposes a heavy transmission burden. The smaller the $\mu_{n}$, the greater the level of semantic compression, which significantly reduces the transmission load. However, semantic extraction introduces additional computational overhead and may decrease task performance on the ES. The minimum value $\mu_{n}^{min}$ is typically determined by task accuracy requirements and user battery capacity. Therefore, selecting $\mu_{n}$ requires a comprehensive evaluation of the current task computational load, channel environment, user battery capacity, and other relevant information.
@@
II-B Transmission Model
The channel gain from UEn to ES via sub-channel $k$ is denoted by $|g_{k,n}|^{2}=|h_{k,n}|^{2}/d_{n}^{-\alpha}$, where $h_{n}\sim\mathcal{C}\mathcal{N}(0,1)$ is Rayleigh fading coefficient, and $d_{n}$ is the distance from UEn to ES, and $\alpha$ is the path loss exponent, and $\mathcal{C}\mathcal{N}(0,1)$ is the complexed Gaussian distribution with zero mean and variance one. The maximum achievable offloading data rate from UEn to ES can be represented as

### v3

# QoE-Driven Multi-Task Offloading for Semantic-Aware Edge Computing Systems

@@ I Introduction
The sixth generation (6G) of mobile communication is demonstrating its revolutionary potential, portending the arrival of an era of ultra-massive connectivity and the Internet of Things (IoT) [1]. The volume of data generated by a vast number of devices is experiencing an explosive surge. Various compute-intensive services, such as holographic communication, extended reality (XR), and autonomous driving continue to evolve, which increasingly burden user equipments (UEs) with heavy computational demands. However, the limited computational power and battery capacity of UEs make it challenging to support a multitude of computationally intensive tasks.
Cloud computing is considered capable of alleviating the computational pressure on UEs to some extent [2]. Offloading tasks to the cloud, which has abundant computing resources, significantly reduces local computational pressure. However, the high latency associated with cloud offloading is intolerable for latency-sensitive tasks. Furthermore, the centralized model of cloud computing poses data privacy and security issues. To address the aforementioned challenges, mobile edge computing (MEC) is considered an effective solution. Edge servers (ES) near the user can provide considerable computing and storage resources, which can be fully utilized by UEs distributed. Compared to cloud computing, MEC significantly reduces latency, while its distributed computing resources ensure user privacy and security.
@@
Nevertheless, when tasks involve substantial data transfers, such as high-definition (HD) video processing [3], virtual reality (VR) applications [4, 5], or real-time data analytics [6, 7], the increased volume of data imposes higher demands on channel quality to meet latency requirements. In such cases, semantic communication (SemCom) can provide an effective solution [8]. SemCom, different from the traditional bit workflow based on Shannon’s information theory [9], has demonstrated great potential for enhancing the efficiency and reliability of data transmission across various task scenarios, including text/audio [10, 11, 12, 13], image [14, 15, 16], video [17, 18, 19], XR [20, 4, 21], and multi-modal [22, 23, 24]. By transmitting a compact semantic representation of task-relevant data, substantial bandwidth can be saved, ensuring high quality of service (QoS).
@@
However, these works lack consideration for task execution performance and comprehensive QoE design. 6G strives to provide stable, reliable, and high-quality personalized services to each user [30]. Preferences often vary among different users and types of services. For instance, in MEC networks, users dealing with latency-sensitive tasks may prefer lower latency, those pursuing task performance may favor better task execution accuracy, while users with limited battery power might opt for task offloading strategies that consume less power. Therefore, a good QoE metric should encompass the preferences of different users. Additionally, current research lacks comprea hensive consideration of the inherent heterogeneity across multi-modal tasks. Since each user may need to handle different tasks, the complexity and evaluation metrics of different tasks vary. For example, bimodal tasks like visual question answering (VQA) typically require more processing time and computational resources compared to text translation, and their evaluation metrics are entirely different.
Based on the above considerations, we design a resource allocation scheme for semantic-aware MEC networks that supports multi-modal tasks. We propose a semantic-aware factor to balance computation, communication, and task execution performance. To the best of our knowledge, this work is the first semantic-aware task offloading framework that jointly incorporates execution latency, execution energy consumption, and task performance into the QoE formulation. Additionally, a fair QoE metric is introduced to evaluate the utility of multi-modal task offloading, with support for user preference settings. The main contributions of this paper are summarized as follows.
  • We propose a novel semantic-aware task offloading system for multi-tasks. The SemCom systems for image classification, text classification, and VQA are first constructed to compress source data into compact semantic representations. The semantic extraction factor is used to characterize the degree of data compression, which affects the task’s execution latency, energy consumption, and execution performance.
@@
The remainder of this paper is organized as follows. Section II introduces the existing literature and related works. Section III describes the semantic-aware transmission and computation model, and defines the unified QoE metric, and formulates the problem as maximizing overall QoE. Section IV introduces solving the QoE maximization problem using the MAPPO algorithm, providing training details. Section V presents relevant experimental details and simulation results. Finally, Section VI outlines our conclusions.
II Related Works
MEC networks have attracted widespread attention from academia and industry [31] due to its ability to enhance the computing capabilities of UEs and improve the QoS. Numerous studies have adopted numerical optimization techniques to derive optimal or near-optimal solutions [32, 33, 34, 35, 36, 37]. In [32], the authors optimized the allocation of communication, computation, and energy resources based on a perturbed Lyapunov algorithm, maximizing system throughput under the stability constraints of task and energy queues. A power minimization method with task buffer stability constraints was proposed in [33], where the Lyapunov optimization algorithm was used to solve computing offloading strategies. The requirements for ultra-reliable and ultra-low latency (URLLC) tasks was considered in [34], applying extreme value theory to minimize user power consumption while satisfying probabilistic queue length constraints. In [35], a multi-branch deep neural network (DNN) model was developed to minimize energy consumption in device–edge collaborative inference, and the trade-off between branch selection and inference accuracy was analytically quantified.
However, the aforementioned model-based approaches are typically computationally intensive, further burdening user devices. Moreover, model-based methods often require multiple iterations to find the optimal offloading strategy, which is intolerable for latency-sensitive tasks and impractical in rapidly changing wireless channel conditions. To address these issues, resource allocation and computation offloading methods based on Deep Reinforcement Learning (DRL) were proposed [38, 39, 40, 41, 42, 43]. The authors of [38] modeled the computation offloading as a multi-stage stochastic mixed-integer nonlinear programming (MINLP) problem and solved it using a combination of Lyapunov optimization and a DRL model, integrating the advantages of both model-based and DRL approaches. In [39], two offloading schemes based on soft actor critic (SAC), i.e., centralized SAC offloading (CSACO) and distributed SAC offloading (DSACO) were proposed for routing and scheduling between adjacent ESs to alleviate congestion in the core network. CSACO collects global states and makes decisions to approximate the global optimum, while DSACO enables each ES to make decisions based on local states, supporting decentralized learning and reducing communication overhead. In [40], the authors designed a multi-agent deep deterministic policy gradient (MADDPG)-based path planning scheme to minimize the total energy consumption of unmanned aerial vehicle (UAV) systems. During training, the agents exploit global information, whereas in actual deployment, they make decisions relying solely on local observations. In [41], a distributed DRL algorithm was proposed to minimize task offloading delay. The authors enhanced the long-term value estimation by integrating long short-term memory (LSTM) and double deep Q-network (DDQN).
The optimization objectives in task offloading can be either single or multiple. Single-objective optimization may aim to minimize offloading delay, energy consumption, or operational cost [44, 45, 46, 47, 48, 49, 50]. In [44], the authors jointly optimized server deployment and user offloading decisions in a wireless backhaul edge network, proposing an efficient genetic algorithm-based approach to minimize the average service delay of users. An aerial hierarchical MEC system was introduced in [45], which supports dynamic task offloading across three tiers of nodes: local devices, unmanned aerial vehicles (UAVs), and high-altitude platforms (HAPs). In [46], the authors constructed a distributed many-to-many matching model that allows task owners and edge nodes to negotiate offloading and payment decisions through multiple rounds of interactive and autonomous bargaining, thereby preserving user privacy. In [47], a novel intelligent framework integrating edge computing with few-shot learning was developed to enable semantic inference based on limited data samples.
Multi-objective optimization in task offloading aims to simultaneously consider multiple performance goals. Typically, several objectives are combined through weighted summation or by treating some as constraints [51, 52, 53, 54, 55, 56, 57, 58]. In [51], a parameterized DQN-based method was proposed to solve the online service placement and resource allocation problem. Two DNNs are employed to estimate the optimal service placement decisions and the corresponding computing resource allocation, respectively. The authors of [52] developed a collaborative inference model that dynamically selects the partition point of deep learning models. By designing an experience-sharing mechanism, neighboring devices exchange Q-values to accelerate the optimization of the policy network. In [53], a three-tier MEC network architecture supporting end–edge–cloud collaborative caching was proposed, aiming to jointly optimize content retrieval delay and handover latency. The authors of [54] were the first to model a task offloading problem with spatio-temporal load awareness, effectively preventing the overload of certain satellites while avoiding resource underuse in others.
In MEC networks, task characteristics such as heterogeneity, dependency, and priority significantly influence offloading decisions [59, 60, 61, 62, 63, 64, 65]. In [59], the authors designed a meta-learning mechanism that enables the algorithm to rapidly adapt to dynamic changes in the edge–cloud environment. In [60], the authors investigated the multi-task offloading problem in vehicular edge computing (VEC), considering dynamic and flexible task topologies, heterogeneous resource demands, and variations in communication and computation capabilities. The authors of [63] modeled interdependent tasks as a directed acyclic graph (DAG) and proposed a priority-based scheduling method derived from task cost, addressing collaborative offloading under multi-task dependencies in satellite IoT systems. In [65], the authors jointly optimized task partitioning ratios and user association strategies in a multi-user, multi-edge-node MEC scenario, supporting the offloading of both independent subtasks and sequentially dependent subtasks.
Semantic encoding leverages powerful feature extractors to drastically reduce required bandwidth and has demonstrated strong potential in text/audio [10, 11, 12, 13], image/video [14, 15, 16, 17, 18, 19], 3D/XR [20, 4, 21], and multi-modal transmission [66, 67]. [10] is one of the earliest implementations of text-based semantic encoding. The authors designed an end-to-end deep learning semantic communication system that jointly optimizes semantic and channel encoding, effectively extracting semantic information from text and enhancing robustness against channel noise. Subsequently, in [10], the authors further proposed a lightweight distributed semantic communication system that significantly reduces model size and complexity through weight pruning and quantization, enabling deployment on resource-constrained IoT devices. In [12], the authors developed a knowledge graph (KG)–based text semantic communication system, in which textual data are represented as “entity–relation–entity” triples. A robust voice semantic transmission scheme was proposed in [13], where the authors adopt a hierarchical design that separates audio compression from digital encoding and transmission, while fully exploiting the contextual prediction capabilities of large language models (LLMs).
Semantic transmission technologies have been widely applied in the field of visual data. In terms of semantic image transmission, a deep learning–based joint source–channel coding (JSCC) method was proposed in [14], , which maps image pixel values to complex-valued channel input symbols, thereby enabling the end-to-end wireless image transmission system. In [68], the authors developed a wireless image semantic communication framework that jointly optimizes system capacity, pixel-level distortion, semantic-level distortion, and perceptual quality. The information bottleneck theory was employed to balance the trade-off between channel bandwidth ratio and transmission robustness. A content-aware image semantic transmission scheme was presented in [69], where images are divided into regions of interest (ROI) and regions of non-interest (RONI) according to downstream tasks. By adjusting quantization levels, the algorithm significantly reduces the amount of encoded data. In [70], a multi-level generative image semantic transmission approach was proposed, in which both textual and visual features are used to guide image denoising to generate realistic images.
In terms of semantic video transmission, a semantic transmission scheme for video conferencing was proposed in [18], which significantly reduces data redundancy by transmitting only the keypoint changes between adjacent frames. In [71], a latent diffusion model (LDM)–based semantic-aware adaptive bitrate streaming framework was designed. By combining LDM with FFmpeg technology, the framework optimizes bandwidth utilization and storage efficiency for real-time video streaming. An ultra-low bitrate semantic video transmission scheme was developed in [72], where only the first frame image and subsequent audio signals are transmitted to drive the lip movements of individuals in the video. A video transmission scheme supporting semantic self-correcting and fine-grained bitrate control is proposed in [73], where video frames are encoded into a semantic codebook space and matched with the current bandwidth through active packet loss.
Due to the massive volume of XR and 3D data, many researchers have explored semantic encoding algorithms for data transmission. In [4], a semantic transmission framework tailored for VR content was developed. The concept of semantic location graph (SLG) was introduced, which significantly reduced data transfer volume by separately transmitting the semantic features of static objects and dynamic objects. In [74], neural radiance fields (NeRF) were combined with nonlinear transform coding for the first time to develop an end-to-end semantic transmission system for free-viewpoint 3D scenes. An interest-aware communication framework for point cloud video (PCV) streaming was proposed in [75]. The authors designed a two-stage ROI selection mechanism that dynamically filters key content based on user motion trajectories and saliency analysis, substantially reducing network and computational resource consumption while maintaining video quality. In [76], an end-to-end deep learning network integrating feature extraction and reconstruction was designed. By transmitting only the key features of point clouds instead of raw data, this approach markedly reduced the bandwidth requirements for real-time point cloud video streaming.
TABLE I: Comparison Between Related Works and Ours
References Semantically Adjustable Multi-Task Energy Efficiency Communication Efficiency Execution Performance Resource Allocation DRL Algorithm Hybrid Action Space
[77]
[78]
[79]
[80]
[81]
[82]
[83]
[84]
[85]
[86]
[87]
[26]
[88]
[29]
[28]
our work
The aforementioned semantic transmission schemes are all based on single-modality. However, in current multi-user and multi-scene environments, data from multiple modalities are often generated. Many studies have therefore sought to exploit multi-modal data to develop more efficient compression methods or design transmission frameworks that support multiple modalities. In [89], a multi-modal image compression framework was proposed, which for the first time utilized the semantic information from textual descriptions as prior knowledge to guide image compression, thereby improving reconstruction quality at extremely low bitrates. In [23], the first unified end-to-end semantic communication system supporting multi-modal data was developed, capable of handling diverse modality tasks within a single model. In [24] a Transformer-based architecture for multi-user semantic communication was introduced, which efficiently extracts semantic information for different tasks, addressing the challenges of semantic extraction in multi-modal and multi-user scenarios. In [90], an intelligent agent system based on a large edge AI model (LAM) was designed for the first time, capable of processing multi-modal data in a unified manner, understanding user intent through natural language, and generating personalized wireless communication strategies.
Semantic encoding technology can further reduce uplink traffic in MEC networks, and many studies have integrated the two approaches [77, 78, 79, 80, 81]. In [87], the authors utilized a shared probabilistic graph for semantic information compression and derived the lower bound of signal estimation error under imperfect CSI. The authors of [86] focused on DNN inference tasks and quantitatively analyzed the nonlinear relationship between task compression ratio and local computation ratio. A semantic-aware MEC network for sustainable next-generation consumer electronic devices was proposed in [85]. By extracting and transmitting the semantic information of tasks using semantic communication, the approach significantly reduces transmission latency and energy consumption. In [84], a multi-base-station, multi-terminal MEC system was designed, which reduces computational overhead by dynamically adjusting the depth of neural network computations. The authors of [83] combined semantic communication with low-Earth orbit (LEO) satellite-based edge clouds, extending the semantic encoder deployment to both ground stations and satellites. In [82], a universal task offloading framework was proposed, integrating semantic communication, non-orthogonal multiple access (NOMA), MEC, and user mobility awareness. This framework introduces adjustable semantic factors for semantic extraction of task data, enabling dynamic trade-offs between local computational overhead and data transmission volume.
Table I presents a comparison between our work and existing works across several key dimensions, including semantic adaptation, multi-task support, energy efficiency, communication efficiency, task execution performance, resource allocation, DRL algorithms, and hybrid action spaces. Leveraging semantic extraction factors, this paper comprehensively considers latency, energy consumption, and task performance while supporting the offloading of diverse heterogeneous tasks in MEC networks.
III System Model and Problem Formulation
@@
We propose a novel semantic-aware multi-task offloading system, as shown in Fig. 1, where UEs employ an orthogonal frequency-division multiple access (OFDMA) scheme. At the access point (AP), $K$ resource blocks are allocated to serve $N$ users at different sub-channels. The set of UEs and sub-channels are defined as $\mathcal{N}=\{1,2,\ldots,N\}$ and $\mathcal{K}=\{1,2,\ldots,K\}$. The length of the UEn’s task queue is denoted as $Q_{n}$. The task currently being executed by the UEn is represented by $q^{\mathcal{M}}_{n}$, where $\mathcal{M}\in\{task_{1},task_{2},\ldots,task_{m}\}$. Hence, each UE may contain up to $m$ different types of tasks. We assume that tasks are atomic [29], i.e., all tasks are either executed locally or offloaded to the ES. Considering the computational burden of semantic extraction and deep reinforcement learning, Graphic Processing Unit (GPU)-based architectures are employed to accelerate both training and inference in our proposed framework. The illustrations of the main notations are summarized in Table II.
III-A Semantic-Aware Model
@@
Traditional MEC offloading methods usually require transmitting all data to the ES. SemCom enables the extraction of task-relevant semantic information, reducing bandwidth consumption while enhancing robustness to channel variations. Based on this, we propose a semantic-aware multi-task offloading framework. Fig. 2 illustrates the process of semantic-aware offloading. While semantic extraction indeed reduces the volume of data transmitted, the associated increase in semantic extraction latency and computational burden must also be comprehensively considered. Hence, we define a semantic extraction factor $\mu_{n}\in\left[\mu_{n}^{min},1\right]$. The semantic extraction module determines the level of data compression. When $\mu_{n}=1$, the UE apply traditional offloading scheme instead of semantic-aware offloading.
The proposed adaptive semantic extraction is inspired by the split computing paradigm in edge intelligence [91, 92], where a deep learning model is partitioned between the UE and the ES. The semantic extraction module executes part of the model locally at the UE, while the remaining computation is handled at the ES using the transmitted intermediate semantic representations. The semantic extraction factor $\mu_{n}$ determines the balance of computational workload between the UE and ES, as well as the size of the transmitted features. While deeper local execution increases computation cost but reduces communication load, excessive compression of intermediate representations may degrade task performance at the server.
To maintain analytical tractability while capturing the essential trade-offs, without loss of generality, we model the semantic extraction overhead, transmission delay, and task performance as monotonic power-law functions of the semantic extraction factor. Such phenomenological modeling has been widely adopted in edge intelligence and communication systems to approximate complex deep learning behaviors [93, 94, 95]. It effectively captures the diminishing returns and non-linear sensitivity of computation and performance with respect to semantic compression. This abstraction enables joint optimization of semantic-aware offloading decisions without loss of generality, while avoiding the prohibitive complexity of explicit neural network-level modeling.
Transmitting all data can achieve better task performance, but it imposes a heavy transmission burden. The smaller the $\mu_{n}$, the greater the level of semantic compression, which significantly reduces the transmission load. However, semantic extraction introduces additional computational overhead and may decrease task performance on the ES. The minimum value $\mu_{n}^{min}$ is typically determined by task accuracy requirements and user battery capacity. Therefore, selecting $\mu_{n}$ requires a comprehensive evaluation of the current task computational load, channel environment, user battery capacity, and other relevant information.
@@
III-B Transmission Model
The channel gain from UEn to ES via sub-channel $k$ is denoted by $|g_{k,n}|^{2}=|h_{k,n}|^{2}/d_{n}^{-\alpha}$, where $h_{n}\sim\mathcal{C}\mathcal{N}(0,1)$ is Rayleigh fading coefficient, and $d_{n}$ is the distance from UEn to ES, and $\alpha$ is the path loss exponent, and $\mathcal{C}\mathcal{N}(0,1)$ is the circularly-symmetric complex Gaussian distribution with zero mean and variance one. The maximum achievable offloading data rate from UEn to ES can be represented as